OpenStudy (anonymous):

Three numbers which are coprime to each other are such that the product of first two is 551 and that of the last two is 1073. Then determine the sum of three numbers.?

OpenStudy (ganpat):

are the answers 19, 29 & 37 ??

OpenStudy (anonymous):

Yes right..

OpenStudy (ganpat):

any specific formula ?@waterineyes ??

OpenStudy (anonymous):

xy = 551 yz = 1073 Divide them: \[\frac{x}{z} = \frac{551}{1073} \implies \frac{19}{37}\]

OpenStudy (ganpat):

wow.. i almost tried the same :D.. anyways, thnx :)

OpenStudy (anonymous):

29 is the factor which is diving both..

OpenStudy (anonymous):

thanks